The dimensions of the official NBA basketball court are 94 feet and 50 feet. How many basketball courts can you fit in the Pentagon if the area is 14 hectare ?
Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

court area = (94)(50) = 4700 ft^2

Pentagon area = 14 hectares = 1,506,946 ft^2

# of courts = 1,506,946/4700 = 320.6 courts

Answer by ikleyn(52879)

.

The answer by @CubeyThePenguine would be correct, had you have a possibility to deform these official NBA basketball courts

topologically by changing their shape while conserving their areas.


But since you can not do it in reality, that answer is nonsensical.